Solve by completing the square; round to the nearest hundredth if necessary.
x^2 -3x = 4

Thank you for your help.

Answer by Boreal(15235)

x^2 -3x = 4
x^2-3x+(9/4)=4+(9/4)=(25/4)
(x+3/2)^2=(25/4)
take the square root of both sides
(x-(3/2))=+/-(5/2)
x=(3/2)+(5/2)=4
x=(3/2)-(5/2)=-1
Check 1-(-3)=4
and 16-12=4
and x^2-3x-4=0
(x-4)(x+1)=0; x=4,-1